OpenStudy (anonymous):

x-2/3^7-2/3^8-2/3^9=1/3^9 then x=

OpenStudy (anonymous):

can you clean that up, and use equation (repost)

OpenStudy (aravindg):

well you can see that on left hand side (2/3 )^7 is common to all terms except x ..so take that common outside

OpenStudy (aravindg):

*-(2/3)^7

OpenStudy (aravindg):

can you do it ?

OpenStudy (anonymous):

nooo, It is not (2/3)^7, but 2/3^7 , 2/(3)^7

OpenStudy (aravindg):

still you can use the same logic take 2/(3&7)

OpenStudy (anonymous):

can you use the equation below and rewrite it neatly?

OpenStudy (aravindg):

2/3^7

OpenStudy (anonymous):

\[x-(\frac{ 2 }{ 3 })^7-(\frac{ 2 }{ 3 })^8-(\frac{ 2 }{ 3 })^9=(\frac{ 1 }{ 3 })^9\]\[x-(\frac{ 2 }{ 3 })^{24}=(\frac{ 1 }{ 3 })^9\]\[x=(\frac{ 1 }{ 3 })^9+(\frac{ 2 }{ 3 })^{24}\]\[x=\frac{ 1+2^{24} }{ 3^{24 }}\approx5.94x10^-5\]

OpenStudy (anonymous):

\[x-\frac{ 2 }{ 3^7 }\] guy all fractions are like this , got it?

OpenStudy (anonymous):

\[x-\frac{ 2 }{ 3^7 }-\frac{ 2 }{ 3^8 }-\frac{ 2 }{ 3^9 }=\frac{ 1 }{ 3^9 }, then x= ??\]

OpenStudy (aravindg):